At the beginning of the school year, the mean score of a group of 24 students in an educational achievement test in reading was 45 with a standard deviation of 6. At the end of the school year, the mean score on an equivalent form of the same test was 50 with a standard deviation of 5. Has the class made significant progress in reading during the year? Use alpha equal to 0.05.


1
Expert's answer
2022-01-05T15:36:48-0500

"H_0:\mu_1=\mu_2" , mean scores at the beginning of year and at the end of year are equal

"H_a:\mu_2>\mu_1" , mean score at the end of year is more than at the beginning of year (the class has made significant progress)


Paired T-Test:

"t=\frac{\mu_2-\mu_1}{s_d/\sqrt n}"


where standard deviation of the difference:

"s_d=\sqrt{s_1^2+s_2^2}=\sqrt{6^2+5^2}=7.81"


"t=\frac{50-45}{7.81/\sqrt { 24}}=3.136"


"df=n-1=23"

critical value:

"t_{crit}=1.714"


Since "t>t_{crit}" we reject null hypothesis. The class has made significant progress.